### v1

@@ Appendix A More Related Works
For random ER graphs with number of nodes $n$ and average degree $d$, the upper bound of MIS size is $\frac{2n\ln d}{d}$ for asymptotically large $n$ and $d$(Coja-Oghlan & Efthymiou, 2015). Grimmett & McDiarmid (1975) proves that the simplest Ran-Greedy can achieve half-optimal at $\frac{n\ln d}{d}$ on random ER graphs. Despite that, there is no known existing polynomial algorithm which can reach MIS size of $(1+{\epsilon})\frac{n\ln d}{d}$ for any constant ${\epsilon}$ for asymptotically large $n$ and $d$(Coja-Oghlan & Efthymiou, 2015). Coja-Oghlan & Efthymiou (2015) also suggests that the reason is likely independent sets of size larger than $(1+{\epsilon})\frac{n\ln d}{d}$ forms an intricately ragged landscape, where local algorithms will stuck. Gamarnik & Sudan (2014); Rahman & Virág (2017) proves that for local algorithms (which is defined to only use information from a constant neighborhood of a node to decide whether the node is in the independent set) are at most half optimal for independent set on random $d$-regular graphs. Gamarnik & Sudan (2014) suggests this is due to a property of the MIS problem, which they denote as the Overlap Gap Property (OGP) (Gamarnik, 2021). Gamarnik (2023) suggests that graph neural networks (GNNs) are also a type of local algorithms and thus being limited by OGP. While most AI-inspired MIS algorithms use GNN, the proof only applies to algorithms use GNNs as the only component to find solutions like (Schuetz et al., 2022), so it may not apply directly to more complicated algorithms like those tested in our paper. Yet, it may still suggests a reason why GNN-based algorithms (including most AI-inspired algorithms) cannot outperform classical heuristics like KaMIS.
In addition, Barbier et al. (2013) provides a conjectured tighter upper bound than $\frac{2n\ln d}{d}$ for $d$-regular graphs using the hard-core model in physics. Ding et al. (2016) proves a similar tighter upper bound for $d$-regular graphs. Wormald (2003) gives average-case performance for Deg-Greedy on $d$-regular graphs.
@@
Over the past few decades, significant progress has been made in tackling NP-hard combinatorial optimization (CO) problems by developing approximation algorithms and heuristic methods. Approximation algorithms provide provable guarantees on solution quality and have led to groundbreaking results for classical problems, such as the Maximum Independent Set (MIS), Traveling Salesperson Problem (TSP), and Maximum Cut (Boppana & Halldórsson, 1992; Laporte, 1992; Goemans & Williamson, 1995).
As we mentioned in Section 5, there are various existing heuristics for MIS. For exmaple, reduction techniques reduce the graph into smaller instances. Akiba & Iwata (2016) and Xiao & Nagamochi (2013) have shown a variety of reduction techniques which work well for MIS problem. Butenko et al. (2002); Bourgeois et al. (2012) uses reduction techniques to develop efficient exact algorithms for MIS. Local search improves an existing independent set by removing a small number of nodes and insert other eligible nodes. Andrade et al. (2012) gives an efficient local search algorithm for MIS and has been used as a subprocess for several MIS solvers. Evolutionary algorithms combine several existing solutions into a new solution. Examples include Back & Khuri (1994); Borisovsky & Zavolovskaya (2003); Lamm et al. (2015). KaMIS (Lamm et al., 2017; Dahlum et al., 2016) was developed based on many of these techniques above.
In addition, the MIS problem can also be relaxed into semi-definite programming (SDP), which leads to several approximation algorithms (Halperin, 2002; Bansal, 2014).
@@
In addition to Section 5, we note that Sun et al. (2022); Sanokowski et al. (2023; 2024) use annealing techniques. Sun & Yang (2023) was developed based on Qiu et al. (2022), and Sanokowski et al. (2024) was based on Sanokowski et al. (2023). Sanokowski et al. (2024) can also be considered as an extension of Karalias & Loukas (2020).
Besides MIS, Maximum Cut, and TSP, people also considered to use AI-Inspired methods for other combinatorial optimization problems, including Vehicle Routing Problems (including TSP) (Kool et al., 2018; Chen & Tian, 2019; Delarue et al., 2020; Li et al., 2021; Zheng et al., 2021; Ye et al., 2024), Job Scheduling Problems (Lin et al., 2019; Baer et al., 2019; Zhang et al., 2020; Ye et al., 2024), Boolean Satisfiability (SAT) (Amizadeh et al., 2019; You et al., 2019; Kurin et al., 2020; Li & Si, 2022; Li et al., 2023), and Casual Discovery (Zheng et al., 2018; Zhu et al., 2020; Sanchez et al., 2022).
@@ B.2 Datasets
ER graphs (Erdös & Rényi, 1959) have 2 parameters: the number of nodes $n$ and the average degree $d$. For graphs with given $(n,d)$ We generate them by choosing $M=\frac{nd}{2}$ edges uniformly at random between $n$ nodes. This is the Erdős-Renyí’s $G(n,M)$ model.
@@
BA graphs (Albert & Barabási, 2002) also have 2 parameters: the number of nodes $n$ and the generation parameter $m$, with $n>m$. For a given $(n,m)$, the generation process initializes with $m$ nodes, and then add 1 node at each step. When adding a new node, $m$ neighbors of the new node are sampled from the existing nodes, with the probability of the current degree of the nodes. The average degree of BA graph with given $(n,m)$ can be computed as $d=2m-\frac{m}{n}-\frac{m^{2}}{n}\approx 2m$.
@@
Since we need at least $4100$ graphs to train and test our algorithms and the graphs should not be too small, it is difficult to find such datasets. Fortunately, Morris et al. (2020) provides a website www.graphlearning.io which includes many graph datasets prepared by them or collected from other works. Although most of the datasets are still too small or having too small graphs, we are able to find 2 datasets: REDDIT-MULTI-5K and COLLA, both from Yanardag & Vishwanathan (2015). REDDIT-MULTI-5K has $4999$ graphs with average nodes $508.52$ and average edges $594.87$, so the graphs are generally very sparse. COLLA has $5000$ graphs with average nodes $74.49$ and $2457.78$, so they are denser but smaller graphs.
We were not able to find real-world datasets with enough size of more larger and denser graphs, which are generally more difficult for MIS algorithms. The dataset DIMACS used in Böther et al. (2022) contains such graphs but they only have $37$ graphs which is not enough for training.
@@ B.4 Classical CPU-based algorithms
We did not include results of Ran-Greedy in the main paper because its performance is significantly lower than all other algorithms. It can be considered as a baseline and has theoretical significance, since it has provable guarantee for random graphs (Grimmett & McDiarmid, 1975).
@@
KaMIS (Karlsruhe Maximum Independent Sets) (Lamm et al., 2017) (https://karlsruhemis.github.io/) is the state-of-art heuristic solver for MIS and has been used as a baseline in many previous works. It provides 2 algorithms for the MIS problem: ReduMIS (Lamm et al., 2017) and OnlineMIS (Dahlum et al., 2016).
@@ B.5 GPU-accelerated non-learning algorithms
iSCO (improved Sampling for Combinatorial Optimization) (Sun et al., 2023) is a GPU-accelerated sampling-based method. It does not require learning. According to (Sun et al., 2023), the main benefit is that it can process a large batch of graphs in parallel thus improve efficiency. While processing a small number of graphs like in our case (8 test graphs), it still requires significant time, often longer than ReduMIS.
We use the code from the codebase of DISCS (Goshvadi et al., 2024), which is a follow-up paper for iSCO, as iSCO did not provide the codebase. We use 1 80GB $A100$ GPU to run iSCO with all test graphs together. The time limit is 96hrs, and the actual time it requires is shorter. It fail to run graphs of size $(n=3000,1000)$, $(n=10000,d=100)$, and larger because they require larger than 80GB memory. The code does not support multi-GPU.
@@
PCQO (Parallelized Clique-Informed Quadratic Optimization) (Alkhouri et al., 2024) is a GPU-accelerated gradient-based optimization algorithm, which directly optimize on the quadratic loss function of the MIS problem. The loss function for a graph $G=(V,E)$ and solution vector ${\bm{x}}$ is
@@
Although Alkhouri et al. (2024) reported that using the third term in the loss function (i.e. nonzero $\gamma^{\prime}$) can lead to better answers comparing to only using 2-term loss function (set $\gamma^{\prime}=0$), we found that a nonzero $\gamma^{\prime}$ makes the algorithm more sensitive to hyperparameters and difficult to perform hyperparameter tuning for our benchmark with many datasets. Therefore, we only use $2$-term loss function by setting $\gamma^{\prime}=0$.
In the grid search for hyperparameters, we use learning rate $0.001$ (default), momentum $\{0.5,0.9\}$, $\gamma\in\{4,20,100,500,2500\}$, and report the results obtained from the best hyperparameters. However, for $n\geq 30000$, the grid search within this domain does not provide solution with reasonable size (worse than Ran-Greedy), so we do not report results for larger graphs. Despite that, there may exist a better set of hyperparamters which could make this algorithm perform well on larger graphs.
@@ B.6 Learning-based algorithms
LwD (Learning what to Defer) (Ahn et al., 2020) is a reinforcement learning based algorithm which requires training data to learn the policy. It models the MIS problem as a Markov Decision Process (MDP), where in each step it selects some (possibly 00, $1$, or multiple) nodes to add into the independent set. It is a non-backtracking algorithm as the added nodes are never taken out. Böther et al. (2022) also included this algorithm in their benchmark.
We use the code from Böther et al. (2022) since it provides better functionality than the original codebase. We use the default setting provided by Böther et al. (2022) in their MIS Benchmark codebase (https://github.com/MaxiBoether/mis-benchmark-framework), but change the number of samples to $20$ (default is $10$) for test sampling, in order to match the best-of-20 sampling in our benchmark.
@@
LTFT (Let the Flows Tell) (Zhang et al., 2023) similar to LwD, also model the MIS problem as a MDP and non-backtracking. The difference is that it only choose $1$ node at each step, making it more similar to Deg-Greedy. The node chosen at each step is chosen by a GFlowNet (Bengio et al., 2021a), which is trained by in-distribution training data.
We use the default setting provide by Zhang et al. (2023) for training with $20$ epochs. By default setting, it has best-of-20 sampling and report the best solution found.
@@
DIFUSCO (Diffusion Solvers for Combinatorial Optimization) (Sun & Yang, 2023) trains a diffusion model using supervised learning to produce a solution for the MIS. The diffusion model provides an entire solution so it is a one-shot algorithm.
@@
We use the default setting in (Sun & Yang, 2023) except that we use $50$ diffusion steps throughout training and testing, and $20$ samples for testing to be aligned with best-of-20 sampling in other methods. We train the model for $50$ epochs (default) for each dataset.
@@
Although Sun & Yang (2023) suggested that DIFUSCO has some generalization ability. We found the performance degrade significantly for out-of-distribution trained models (specifically trained on smaller graphs with the same average degree but test on larger graphs), we did not report the results of larger graphs where the in-distribution training cannot finish.
@@
DiffUCO (Diffusion for Unsupervised Combinatorial Optimization) (Sanokowski et al., 2024) is also a diffusion model based algorithm but unlike DIFUSCO it uses unsupervised learning. The diffusion model is trained to sample the solution of low energy state. It also provides an entire solution and is also a one-shot algorithm.
We use the default setting in Sanokowski et al. (2024) for RB-large MIS task (in their Appendix C.5). During testing, we use conditional expectation with $20$ samples to align with best-of-20 sampling in other algorithms. The code supports multi-GPU. We use 4 80GB $A100$ GPU to train for each datasets with time limit 96hrs.
@@ B.7 Local search
Local search is a method to improve a given independent set. It can be used as a post-processing technique, or be used as sub-procedures in more complicated algorithms like KaMIS. Andrade et al. (2012) provides an efficient local search algorithm. Part of it is to find $2$-improvement, which is the part used as sub-procedure in KaMIS (Lamm et al., 2017; Dahlum et al., 2016).
The local search algorithm for $2$-improvement for a given independent set $I$ is as follows. This algorithm process every vertex $x\in I$ in turn. First, it temporarily removes $x$ from $I$, creating a new set $S$. We call a vertex a free vertex of $S$ if there is no edge between it and any vertex in $S$. If $S$ has less than two free vertices, stop: there is no 2-improvement involving $x$. Otherwise, for each neighbor $v$ of $x$ that is a free vertex for $S$, insert $v$ into $S$ and check if the new set ($S^{\prime}$) has a free vertex $w$. If it does, inserting $w$ leads to a $2$-improvement; if it does not, remove $v$ from $S^{\prime}$ (thus restoring $S$) and process the next neighbor of $x$. If no improvement is found, reinsert $x$ into $S$ to turn it back to $I$. Every vertex is scanned $O(1)$ times in this algorithm so it can find a $2$-improvement (if there exists) in $O(m)$ time according to Andrade et al. (2012).

### v2

@@ Appendix A More Related Works
For random ER graphs with number of nodes $n$ and average degree $d$, the upper bound of MIS size is $\frac{2n\ln d}{d}$ for asymptotically large $n$ and $d$[51]. [52] proves that the simplest Ran-Greedy can achieve half-optimal at $\frac{n\ln d}{d}$ on random ER graphs. Despite that, there is no known existing polynomial algorithm which can reach MIS size of $(1+{\epsilon})\frac{n\ln d}{d}$ for any constant ${\epsilon}$ for asymptotically large $n$ and $d$[51]. [51] also suggests that the reason is likely independent sets of size larger than $(1+{\epsilon})\frac{n\ln d}{d}$ forms an intricately ragged landscape, where local algorithms will stuck. [49, 53] proves that for local algorithms (which is defined to only use information from a constant neighborhood of a node to decide whether the node is in the independent set) are at most half optimal for independent set on random $d$-regular graphs. [49] suggests this is due to a property of the MIS problem, which they denote as the Overlap Gap Property (OGP) [50]. [25] suggests that graph neural networks (GNNs) are also a type of local algorithms and thus being limited by OGP. While most AI-inspired MIS algorithms use GNN, the proof only applies to algorithms use GNNs as the only component to find solutions like [23], so it may not apply directly to more complicated algorithms like those tested in our paper. Yet, it may still suggests a reason why GNN-based algorithms (including most AI-inspired algorithms) cannot outperform classical heuristics like KaMIS.
In addition, [54] provides a conjectured tighter upper bound than $\frac{2n\ln d}{d}$ for $d$-regular graphs using the hard-core model in physics. [55] proves a similar tighter upper bound for $d$-regular graphs. [56] gives average-case performance for Deg-Greedy on $d$-regular graphs.
@@
Over the past few decades, significant progress has been made in tackling NP-hard combinatorial optimization (CO) problems by developing approximation algorithms and heuristic methods. Approximation algorithms provide provable guarantees on solution quality and have led to groundbreaking results for classical problems, such as the Maximum Independent Set (MIS), Traveling Salesperson Problem (TSP), and Maximum Cut [57, 58, 59].
As we mentioned in Section 2, there are various existing heuristics for MIS. For exmaple, reduction techniques reduce the graph into smaller instances. [32] and [31] have shown a variety of reduction techniques which work well for MIS problem. [30, 60] uses reduction techniques to develop efficient exact algorithms for MIS. Local search improves an existing independent set by removing a small number of nodes and insert other eligible nodes. [33] gives an efficient local search algorithm for MIS and has been used as a subprocess for several MIS solvers. Evolutionary algorithms combine several existing solutions into a new solution. Examples include [34, 35, 36]. KaMIS [28, 37] was developed based on many of these techniques above.
In addition, the MIS problem can also be relaxed into semi-definite programming (SDP), which leads to several approximation algorithms [61, 62].
@@
In addition to Section 2, we note that [43, 41, 16] use annealing techniques. [15] was developed based on [40], and [16] was based on [41]. [16] can also be considered as an extension of [42].
Besides MIS, Maximum Cut, and TSP, people also considered to use AI-Inspired methods for other combinatorial optimization problems, including Vehicle Routing Problems (including TSP) [8, 63, 64, 65, 66, 67], Job Scheduling Problems [68, 69, 70, 67], Boolean Satisfiability (SAT) [71, 72, 73, 74, 75], and Casual Discovery [76, 77, 78].
@@ B.2 Datasets
ER graphs [45] are random graphs where edges are connected uniformly at random (with a fixed probability or number of edges). We vary $2$ parameters for ER graphs, number of nodes $n$ and average degree $d$, by fixing number of edges at $\frac{nd}{2}$. ER graphs are fundamental objects in network sciences [79] and random graph theory [80]. There are also theoretical analysis and conjecture upper bound for MIS on ER graphs [51]. Previous works [24, 7, 15, 21, 18] used it as test graphs for MIS, though without varying parameters as we did.
ER graphs have 2 parameters: the number of nodes $n$ and the average degree $d$. For graphs with given $(n,d)$ We generate them by choosing $M=\frac{nd}{2}$ edges uniformly at random between $n$ nodes. This is the Erdős-Renyí’s $G(n,M)$ model.
@@
Barabási–Albert (BA) graphs [46] are random graphs generated by a probabilistic growth process, whereby new nodes preferentially attach to existing nodes with higher degrees, mimicking real-world networks such as Internet, citation networks, and social networks [46, 81]. For BA graphs, we vary $2$ parameters: number of nodes $n$ and parameter $m$ (not number of edges). The average degree of BA graphs can be approximated as $2m$.
BA graphs also have 2 parameters: the number of nodes $n$ and the generation parameter $m$, with $n>m$. For a given $(n,m)$, the generation process initializes with $m$ nodes, and then add 1 node at each step. When adding a new node, $m$ neighbors of the new node are sampled from the existing nodes, with the probability of the current degree of the nodes. The average degree of BA graph with given $(n,m)$ can be computed as $d=2m-\frac{m}{n}-\frac{m^{2}}{n}\approx 2m$.
@@
RB graphs
RB graphs are derived from Model RB [47], a random constraint satisfaction problem (CSP) model. RB graphs are considered difficult instances for MIS due to their structured randomness and high solution hardness.
We use two datasets from [21] (also used in [16]), RB-small (200–300 nodes) and RB-large (800–1200 nodes), to benchmark learning-based solvers.
@@
We pick REDDIT-MULTI-5K and COLLAB [48] from TUDataset website [82], since they have enough graphs for training and graph sizes not too small. REDDIT-MULTI-5K has $508.52$ average nodes and $594.87$ average edges. They are mostly very sparse graphs. COLLAB has $74.49$ average nodes and 2457.78 average edges. They are mostly small but dense graphs.
Since we need at least $4100$ graphs to train and test our algorithms and the graphs should not be too small, it is difficult to find such datasets. Fortunately, [82] provides a website www.graphlearning.io which includes many graph datasets prepared by them or collected from other works. Although most of the datasets are still too small or having too small graphs, we are able to find 2 datasets: REDDIT-MULTI-5K and COLLA, both from [48]. REDDIT-MULTI-5K has $4999$ graphs with average nodes $508.52$ and average edges $594.87$, so the graphs are generally very sparse. COLLA has $5000$ graphs with average nodes $74.49$ and $2457.78$, so they are denser but smaller graphs.
We were not able to find real-world datasets with enough size of more larger and denser graphs, which are generally more difficult for MIS algorithms. The dataset DIMACS used in [24] contains such graphs but they only have $37$ graphs which is not enough for training.
@@ B.4 Classical CPU-based algorithms
We did not include results of Ran-Greedy in the main paper because its performance is significantly lower than all other algorithms. It can be considered as a baseline and has theoretical significance, since it has provable guarantee for random graphs [52].
@@
KaMIS (Karlsruhe Maximum Independent Sets) [28] (https://karlsruhemis.github.io/) is the state-of-art heuristic solver for MIS and has been used as a baseline in many previous works. It provides 2 algorithms for the MIS problem: ReduMIS [28] and OnlineMIS [37].
@@ B.5 GPU-accelerated non-learning algorithms
iSCO (improved Sampling for Combinatorial Optimization) [17] is a GPU-accelerated sampling-based method. It does not require learning. According to [17], the main benefit is that it can process a large batch of graphs in parallel thus improve efficiency. While processing a small number of graphs like in our case (8 test graphs), it still requires significant time, often longer than ReduMIS.
We use the code from the codebase of DISCS [83], which is a follow-up paper for iSCO, as iSCO did not provide the codebase. We use 1 80GB $A100$ GPU to run iSCO with all test graphs together. The time limit is 96hrs, and the actual time it requires is shorter. It fail to run graphs of size $(n=3000,1000)$, $(n=10000,d=100)$, and larger because they require larger than 80GB memory. The code does not support multi-GPU.
@@
PCQO (Parallelized Clique-Informed Quadratic Optimization) [18] is a GPU-accelerated gradient-based optimization algorithm, which directly optimize on the quadratic loss function of the MIS problem. The loss function for a graph $G=(V,E)$ and solution vector ${\bm{x}}$ is
@@
We use the hyperparamter search range suggested by the authors [18]. We use learning rate $\{0.0001,0.00001,0.000001\}$, momentum $0.9$, $\gamma\in\{200,500,1000,2000,5000\}$, $\gamma^{\prime}=1$, batch size $256$, number of steps $27000$, steps per batch $450$. We then report the results obtained from the best set of hyperparameter. However, for $n\geq 30000$, the grid search within this domain does not provide solution with reasonable size (worse than Ran-Greedy), so we do not report results for larger graphs. Despite that, there may exist a better set of hyperparamters which could make this algorithm perform well on larger graphs. Automatic hyperparameter search could significantly improve the usability of this algorithm.
@@ B.6 Learning-based algorithms
LwD (Learning what to Defer) [7] is a reinforcement learning based algorithm which requires training data to learn the policy. It models the MIS problem as a Markov Decision Process (MDP), where in each step it selects some (possibly 00, $1$, or multiple) nodes to add into the independent set. It is a non-backtracking algorithm as the added nodes are never taken out. [24] also included this algorithm in their benchmark.
We use the code from [24] since it provides better functionality than the original codebase. We use the default setting provided by [24] in their MIS Benchmark codebase (https://github.com/MaxiBoether/mis-benchmark-framework), but change the number of samples to $20$ (default is $10$) for test sampling, in order to match the best-of-20 sampling in our benchmark.
@@
LTFT (Let the Flows Tell) [21] similar to LwD, also model the MIS problem as a MDP and non-backtracking. The difference is that it only choose $1$ node at each step, making it more similar to Deg-Greedy. The node chosen at each step is chosen by a GFlowNet [19], which is trained by in-distribution training data.
We use the default setting provide by [21] for training with $20$ epochs. By default setting, it has best-of-20 sampling and report the best solution found.
@@
DIFUSCO (Diffusion Solvers for Combinatorial Optimization) [15] trains a diffusion model using supervised learning to produce a solution for the MIS. The diffusion model provides an entire solution so it is a one-shot algorithm.
@@
We use the default setting in [15] except that we use $50$ diffusion steps throughout training and testing, and $20$ samples for testing to be aligned with best-of-20 sampling in other methods. We train the model for $50$ epochs (default) for each dataset.
@@
Although [15] suggested that DIFUSCO has some generalization ability. We found the performance degrade significantly for out-of-distribution trained models (specifically trained on smaller graphs with the same average degree but test on larger graphs), we did not report the results of larger graphs where the in-distribution training cannot finish.
@@
DiffUCO (Diffusion for Unsupervised Combinatorial Optimization) [16] is also a diffusion model based algorithm but unlike DIFUSCO it uses unsupervised learning. The diffusion model is trained to sample the solution of low energy state. It also provides an entire solution and is also a one-shot algorithm.
We use the default setting in [16] for RB-large MIS task (in their Appendix C.5). During testing, we use conditional expectation with $20$ samples to align with best-of-20 sampling in other algorithms. The code supports multi-GPU. We use 4 80GB $A100$ GPU to train for each datasets with time limit 96hrs.
@@ B.7 Local search
Local search is a method to improve a given independent set. It can be used as a post-processing technique, or be used as sub-procedures in more complicated algorithms like KaMIS. [33] provides an efficient local search algorithm. Part of it is to find $2$-improvement, which is the part used as sub-procedure in KaMIS [28, 37].
The local search algorithm for $2$-improvement for a given independent set $I$ is as follows. This algorithm process every vertex $x\in I$ in turn. First, it temporarily removes $x$ from $I$, creating a new set $S$. We call a vertex a free vertex of $S$ if there is no edge between it and any vertex in $S$. If $S$ has less than two free vertices, stop: there is no 2-improvement involving $x$. Otherwise, for each neighbor $v$ of $x$ that is a free vertex for $S$, insert $v$ into $S$ and check if the new set ($S^{\prime}$) has a free vertex $w$. If it does, inserting $w$ leads to a $2$-improvement; if it does not, remove $v$ from $S^{\prime}$ (thus restoring $S$) and process the next neighbor of $x$. If no improvement is found, reinsert $x$ into $S$ to turn it back to $I$. Every vertex is scanned $O(1)$ times in this algorithm so it can find a $2$-improvement (if there exists) in $O(m)$ time according to [33].